OpenStudy (chunkymonkay):

what is the solution of the system? y=1/3x y=x-4

OpenStudy (chunkymonkay):

@Hero

OpenStudy (crashonce):

solve simultaneous equations because both are equal to y, solve 1/3x = x-4

OpenStudy (chunkymonkay):

how would i solve that?

OpenStudy (crashonce):

1/3x = x-4, move all x's to one side.

OpenStudy (chunkymonkay):

move all x's to one side so would it look like this 1/3-4 x x i dont know im really confused

OpenStudy (crashonce):

ok so move 1/3x and the x on the right hand side to one side, then leave the rest on one side

OpenStudy (chunkymonkay):

1/3x x -4

OpenStudy (chunkymonkay):

uhm im not sure that we are doing this correctly :/

OpenStudy (chunkymonkay):

i think we must solve for both equations. im not sure if i am correct either but i tryed by myself and here is what i got.

OpenStudy (chunkymonkay):